OpenStudy (babynini):

To find the volume of the solid obtained by rotating the region bounded by the given curves about specified line x=y^2, x=1-y^2 ; about x = 3

OpenStudy (babynini):

to find where they intersect \[\sqrt{x}=\sqrt{1-x}\] so x = 1/2

OpenStudy (babynini):

I'm not sure how to find where they intersect on y o.0

OpenStudy (babynini):

@Empty this is the same type of one zep and I were just doing. But I can't :o

OpenStudy (babynini):

 ...something like that xD